Write the equation of the line below in slope-intercept form and standard form:
Is parallel to y=x+2 and passes through the point (1,-4)

Answer by Alan3354(69443)

Write the equation of the line below in slope-intercept form and standard form:
Is parallel to y=x+2 and passes through the point (1,-4)
----------------
Parallel means it has the same slope. The eqn is in slope-intercept form,
y = mx + b, so the slope of y=x+2 is +1.
------------
Then,
y-y1 = m*(x-x1)
y-(-4) = 1*(x - 1)
y+4 = x-1
y = x - 5 Slope-intercept form,
x - y = 5 Standard form